40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 13 21:04:37 2013, Jenne, Update, Green Locking, PSL green cleaned up 
    Reply  Thu Mar 14 04:20:54 2013, Jenne, Update, Green Locking, Ybeat attempt 
       Reply  Thu Mar 14 20:22:33 2013, Jenne, Update, Green Locking, Xbeat attempt 
    Reply  Thu Mar 14 13:38:29 2013, Jenne, Update, Green Locking, PSL green cleaned up 
Message ID: 8291     Entry time: Thu Mar 14 04:20:54 2013     In reply to: 8290     Reply to this: 8297
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Ybeat attempt 

I dedicated my evening to trying to get the Ygreen beatnote (the idea being to then get the Xgreen beatnote).

First up was tweaking up the green alignment.  Per Yuta's suggestion, elog 8283, I increased the refl PD gain by 2 clicks (20dB) to keep the lock super stable while improving the alignment.  After I finished, I turned it back to its nominal value.  I discovered that I need lenses in front of the DC PD (for Ygreen, and I'm sure Xgreen will be the same).  The beam is just barely taking up the whole 2mm diode, so beam jitter translates directly to DC power change measured by the diode.  I ended up going just by the green transmission camera for the night, and achieved 225uW of Ygreen on the PSL table.  This was ~2,000 counts, but some of the beam is always falling off the diode, so my actual counts value should be higher after installing a lens. 

I then opened up the PSL green shutter, which is controlled by the button labeled "SPS" on the shutter screen - I will fix the label during some coffee break tomorrow.  Using my convenient new PSL green setup, removing the DC PD allows the beam to reflect all the way to the fuse box on the wall, so you can check beam overlap between the PSL green and the arm green at a range of distances.  I did this for Ygreen, and overlapped the Ygreen and PSL green. 

I checked the situation of the beat cabling, since Jamie has the beatbox out for whitening filter modifications tonight.  In order to get some signal into the control room, I connected the output of the BBPD amplifier (mounted on the front of the 1X2 rack) directly to the cable that goes to the control room.  (As part of my cleanup, I put all the cables back the way I found them, so that Jamie can hook everything back up like normal when he finishes the beatbox.) 

I then started watching the signal on the 8591E analyzer, but didn't magically see a peak (one can always hope....).

I decided that I should put the offset in the Y AUX laser slow servo back to the value that we had been using for a long time, ~29,000 counts.  This is where things started going south.  After letting that go for a minute or two, I thought to go check the actual temperature of the laser head.  The "T+" temperature on the controller read something like 42C, but the voltmeter which reads a voltage proportional to the temp (10C/V) was reading 5.6V.  I immediately turned off the offset, but it's going to take a while for it to cool down, so I'll come back in the morning.  I want the AUX laser to be something like 34C, so I just have to wait.  Ooops.

Still to do (for the short-term FPMI):

* Find Y beatnote.

* Align Xgreen to the arm - it's still off in pitch.

* Align Xgreen and PSL green to be overlapped, hitting the BBPD.

* Find the X beatnote.

* Reinstall the beatbox.

* Use ALS to stabilize both arms' lengths.

* Lock MICH with AS.

* Look at the noise spectrum of AS - is there more noise than we expect (Yuta and Koji saw extra noise last summer), and if so, where does it come from?  Yuta calculated (elog 6931) that the noise is much more than expected from just residual arm motion.

* Write a talk.

ELOG V3.1.3-